OpenStudy (anonymous):

Solve the logarithmic equation. Express your solutions in exact form only. Please show all of your work. log 2 (x+8) - log2 (x-4) + log2 (x-5)=3

OpenStudy (anonymous):

\[\log_{2}\frac{(x+8)(x-5)}{x-4}=3\] \[\frac{(x+8)(x-5)}{x-4}=2^3\]

OpenStudy (anonymous):

can you solve for x?

OpenStudy (anonymous):

Idk how to do this thats why I need the answer.

OpenStudy (anonymous):

well i've removed the logarithms, now its just a simple equation

OpenStudy (anonymous):

multiply the right side by (x-4), then subtract the right side from the left side and solve the quadratic